Draw pictures to represent the following: i) 3/2 x 2/3 ii) 0.25 + 0.01 + 0.1 iii) 7/8 < 12/13 iv) 0.1 > 0.08 v) 1/3 of 2/5
1
Expert's answer
2020-04-09T15:07:55-0400

v) "1\/3 of 2\/5 =1\/3*(2\/5)=2\/15"



iv) 0.1>0.08




iii) 7/8 < 12/13 , "7\/8=0.875" and "12\/13=0.923"




ii) "0.25 + 0.01 + 0.1 = 0.26+0.1 = 0.36"




i) "(3\/2) *(2\/3) =6\/6=1"